Circular No. 7686

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2001 Q2 Vance Avery Petriew, Regina, SK, reports his visual discovery of a

comet during a star party at Cyprus Hills, Saskatchewan, as shown below. The object’s presence was also confirmed visually by R. Huziak (0.25-m reflector) and P. Campbell (0.32-m reflector) at Cyprus Hills. The observations by A. Hale were made following a request by the Central Bureau.

2001 UT  α_{2000}  δ_{2000}  m_{1} Observer
Aug. 18.42  5^{h}31.^{m}9  +28^{o}08′  11.0 Petriew
19.45502 5 37 59.^{s}10 +27 47 07′′.8 13.2 Hale
19.46173 5 38 01.45 +27 46 58.4 13.0  _{”}
19.47459 5 38 05.77 +27 46 45.4 13.4  _{”}
19.48078 5 38 07.88 +27 46 36.9 13.4  _{”}

V. A. Petriew (Cyprus Hills, SK). Round coma of diameter 3′ with condensed nucleus and no tail. 0.51-m f/5 reflector at 80×. Motion ∼ 2′ to the southeast over an hour. Magnitude approximate.

A. Hale (Cloudcroft, NM). 0.20-m Schmidt-Cassegrain reflector + CCD. Visual observations with a 0.41-m reflector on Aug. 19.47 showed a coma diameter of 2′.5 and m_{1} = 11.0.

NOVA CYGNI 2001 No. 2 A. Nakamura, Kume, Ehime, Japan, reports the discovery by Akihiko

Tago (Tsuyama, Okayama) of an apparent nova (mag 8.8) on two T-Max 400 films taken on Aug. 18.599 and 18.603 UT with an unfiltered 105mm f/4 camera lens. Tago adds that nothing is visible at this position on his films taken on Aug. 15.62, 16.64, and 17.64 (limiting mag 12.0). H. Abe (Yatsuka, Shimane) measured the following accurate position and magnitude from five unfiltered CCD frames taken with a 0.26-m f/6 reflector on Aug. 19.6 UT: α = 21^{h}03^{m}02^{s}.00, δ = +48^{o}45′52′′.9 (equinox 2000.0), mag 6.6.

K. Ayani, Bisei Astronomical Observatory (BAO), reports that a spectrum (range 460–680 nm; resolution 0.6 nm at Hα), obtained on Aug. 19.57 UT with BAO 1.01-m telescope, shows emission lines of Hα and Hβ, both showing deep P-Cyg absorption: “The FWHM of the Hα emission profile is ∼ 1400 km/s, and the minimum of Hα P-Cyg absorption has a blueshift of 1700 km/s with respect to the emission peak. These features suggest that this is a nova.”

2001 August 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green